OpenStudy (anonymous):

How do you work out these?? a) cos(90°-x) b) tan(180°+x) c) sin(x+180°)

OpenStudy (anonymous):

cos(A-B)=cosA cosB+sinA sinB

OpenStudy (anonymous):

and cos90=0 sin90=1

OpenStudy (anonymous):

tan (A + B) = (tan A + tan B)/(1 - (tan A)(tan B)) and tan180=0

OpenStudy (anonymous):

sin (A + B) = (sin A)(cos B) + (cos A)(sin B) and sin180=0 cos180=-1

OpenStudy (anonymous):

I don't really understand what you're doing, but the question says simplify, if that helps.

OpenStudy (anonymous):

yes apply the rule

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

@esamalaa how?

OpenStudy (anonymous):

in a) use the rule cos(A-B)=cosA cosB+sinA sinB where in ur problem A=90 B=x

OpenStudy (anonymous):

i haven't learnt that rule, is there another way?

OpenStudy (anonymous):

btw the answer for a) is sinx

zepdrix (zepdrix):

Bahhh I'm trying to think of a good way to explain the first one >.<

zepdrix (zepdrix):

The second one isn't too bad.

zepdrix (zepdrix):

Tangent is periodic in pi ( or 180 degrees). So if you add 180 degrees to the angle, it doesn't change tangent at all. \[\Large\bf\sf \tan(180^o+x)\quad=\quad \tan(x)\]

zepdrix (zepdrix):

Too weird? :U Confusing? I could maybe draw a graph if it will help..

OpenStudy (anonymous):

yes please i don't think i've learnt that before

zepdrix (zepdrix):

Tangent is kind of a weird looking graph. Hopefully this doesn't make it more confusing lol.